# DrunkAgent: Stealthy Memory Corruption in LLM-Powered Recommender Agents

### v2

@@ C.2.2 Ablation Studies
We remove important components of each module to conduct ablation studies of DrunkAgent. First, we consider remove the surrogate module (denoted as $\mathcal{S}_{u}$). However, if we remove it, the setting will transfer from the black-box to the white-box that is impractical and beyond the scope of this paper. Moreover, due to gpt-3.5-turbo used for victim models is closed-source and only can be called via APIs, we cannot obtain logits to optimize the adversarial triggers like Llama-3 used in the surrogate model does. Hence, we maintain the surrogate model to ensure more practical settings. We then remove the greedy optimization algorithm (denoted as $\mathcal{G}_{r}$) in the generation module to observe how the performance changes. As illustrated in Table 5, the attack performance moderately degrades after removing $\mathcal{G}_{r}$, which demonstrates the effectiveness of the optimization algorithm. The lack of dramatic performance degradation is probably due to the effectiveness of our initialization. Finally, we mask the strategy module (denoted as $\mathcal{S}_{t}$). From the table, we can find that the attack performance decreases by a large margin. This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines). Our reported results are built upon AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
C.2.3 Parameter Sensitivity
@@
From Table 5, we can find that the strategy module contributes more than the optimization algorithm on DrunkAgent’s attack performance. As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper [5]. As illustrated in Table 6, the success rates of getting the target agents drunk are maintained at 100% regardless of the number of iterations. For the other attack parameters, based on our observations, our DrunkAgent is insensitive to them.
C.2.4 Attack Transferability to Victim Agentic RSs with Powerful LLM backbones
To further evaluate the transferability of DrunkAgent, we improve the LLM backbones of victim agentic RSs with more powerful gpt-4-turbo. Due to such API calls are quite expensive, we evaluate DrunkAgent on a representative real-world dataset. To ensure a sufficiently challenging evaluation, we choose the largest dataset: CD & Vinyl. Fig. 7 shows HR@1 (i.e., NDCG@1) of the benign status and DrunkAgent against the three victim models on the dataset, which demonstrates that DrunkAgent is still transferable even if the agentic RSs become powerful. We leave experiments on additional datasets and more powerful agentic RSs for future work.
Figure 7: Attack Transferability on Powerful Victim Agentic RSs.
@@ C.2.4 Attack Transferability to Victim Agentic RSs with Powerful LLM backbones
Figure 8: Attack Transferability. HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of the Benign status and DrunkAgent against various black-box victim LLM-powered agent-based RSs on the real-world Yelp dataset.
C.2.5 Attack Generalizability across Domains
@@
C.2.6 Attack Interpretability
We focus on why adversarial attacks exhibit both transferability and stealthiness. The transferability arises because adversarial texts are optimized to carry rich as well as general features and explicit attack objectives (e.g., ‘top-level consideration’, ‘a must-have’, ‘prime choice’) to mislead the decisions of the victim models, making them effective across models that share similar language understanding capabilities. The stealthiness is attributed to the fact that these texts are optimized to retain meaningful semantics and coherence, and mainly perturb the memories of the target item agents.
C.2.7 Computing Resourses
All experiments were run on a institution high-performance computing cluster equipped with NVIDIA H100 GPUs (each with 96GB VRAM), using CUDA version 12.4 and NVIDIA driver version 550.127.05. Each individual experimental run (from initialization to evaluation) completed in under 2 hours, with peak GPU memory usage remaining below 30GB. These resource requirements apply across all datasets, methods, and target systems evaluated in the paper.
Appendix D Limitations
While our work demonstrates the effectiveness of memory-based attacks against agentic recommender systems in controlled experimental settings, there are several limitations to be acknowledged.
Our evaluation is conducted on publicly available benchmark datasets and simulated agent environments. Although these setups capture key characteristics of real-world agentic systems, they may not fully reflect the complexity and scale of industrial deployments. In particular, factors such as heterogeneous user behaviors, evolving content ecosystems, and industrial-scale memory management pose additional challenges that are beyond the scope of this study. Future work is needed to validate the generalizability and scalability of our approach in production environments.
In addition, our study focuses on demonstrating the attack effectiveness under a black-box threat model, assuming limited access to internal model parameters or memory states. While this setting reflects practical constraints faced by external attackers, a deeper exploration of defense mechanisms and potential mitigation strategies is left for future investigation.
Despite these limitations, we believe our findings offer valuable insights into the security risks posed by memory mechanisms in agentic recommender systems, and lay a foundation for further research on robust and trustworthy recommendation agents.
Appendix E Broader Impact
As agentic recommender systems are emerging as a promising paradigm for next-generation personalization [11, 8], our work contributes positively by proactively identifying potential security vulnerabilities in these systems. While such vulnerabilities could lead to potential negative societal impacts if misused, our research serves as a foundational step toward building more robust and trustworthy agentic RSs, enabling early risk awareness and informing effective mitigation strategies, thereby helping to minimize negative societal impacts before real-world deployment. Overall, we believe the positive impact of promoting secure and responsible development in this emerging area outweighs the potential risks.

### v3

@@ A.2.1. Ablation Studies
We remove important components of each module to conduct ablation studies of DrunkAgent. First, we consider remove the surrogate module (denoted as $\mathcal{S}_{u}$). However, if we remove it, the setting will transfer from the black-box to the white-box that is impractical and beyond the scope of this paper. We thus choose to maintain the surrogate model to ensure more practical settings. We then remove the greedy optimization algorithm (denoted as $\mathcal{G}_{r}$) in the generation module to observe how the performance changes. As illustrated in Table 5, the attack performance moderately degrades after removing $\mathcal{G}_{r}$, which demonstrates the effectiveness of the optimization algorithm. The lack of dramatic performance degradation is probably due to the effectiveness of our initialization. Finally, we mask the strategy module (denoted as $\mathcal{S}_{t}$). From the table, we can find that the attack performance decreases by a large margin. This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines, which are static and one-off). The reported results are built upon representative AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
A.2.2. Parameter Sensitivity
@@
From Table 5, we can find that the strategy module contributes more than the optimization algorithm on DrunkAgent’s attack performance. As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper (Zhang et al., 2024b). As illustrated in Table 6, the success rates of getting the target agents drunk are maintained at 100% regardless of the number of iterations. For the other attack parameters, based on our observations, our DrunkAgent is insensitive to them.
@@
Figure 7. Attack Transferability. HRs and NDCGs of the Benign status and DrunkAgent against various black-box victim LLM-powered agent-based RSs on the real-world Yelp dataset.
A.2.3. Attack Generalizability across Domains
@@
A.2.4. Attack Interpretability
We focus on why adversarial attacks exhibit both transferability and stealthiness. We attribute the transferability of our adversarial texts to the use of a surrogate model that simulates the victim agent’s behaviors, such as memory updates and recommendations. By optimizing on this surrogate, the texts are equipped with transferable update masking strategies and are encapsulated with generalizable attack objectives (e.g., ‘top-level consideration’, ‘a must-have’, ‘prime choice’) that can effectively mislead similar models. Their stealthiness arises from maintaining semantic meaningfulness and coherence, which conceals manipulative intent and subtly disrupts prompt parsing.